OpenStudy (anonymous):

express the following in simplest radical form 4^1/2 x^1/6 y^2/3

OpenStudy (anonymous):

\[\sqrt{4}\sqrt[6]{x}\sqrt[3^{}]{y^{2}}\]

OpenStudy (anonymous):

 your first wasnt right